OpenStudy (anonymous):

Prove the trigonometric identity: (csc^2x)/(cotx)=cscxsecx

OpenStudy (lalaly):

csc²(x) / cot(x) = csc(x) * csc(x)/cot(x) = csc(x) * [1/sin(x)]/[cos(x)/sin(x)] = csx(x) * [1/sin(x)]*[sin(x)/cos(x)] = csc(x) * [1/cos(x)] = csc(x) sec(x)

OpenStudy (anonymous):

you are pro at trig identities O.o

OpenStudy (lalaly):

heheh thanxx (A)

OpenStudy (anonymous):

Pretty much, you're is kinda saving my life currently. I greatly appreciate it. (:

OpenStudy (lalaly):

:D glad i cd help

OpenStudy (anonymous):

Quick question, how does the /(cosx/sinx) get to *(sinx/cosx) ??

OpenStudy (lalaly):

sory i was away when dividin 1/sinx by cosx/sinx its tha same as 1/sinx * sinx/cosx \[(1/a)\div b = (1/a)\times1/b\]